OpenStudy (anonymous):

For the equation x2 + 3x + j = 0, find all the values of j such that the equation has two real number solutions. Show your work.

OpenStudy (anonymous):

let j=ab a+b=3 check how many solutions exists

OpenStudy (anonymous):

j=3a-a^2 put a=0,a=1,a=2,a=3

OpenStudy (anonymous):

actually \[b^2-4ac>0 \implies 9-4j>0 \] \[j<9/4\]

OpenStudy (anonymous):

Solve the following for x:\[x^2+3x+j=0\]\[x=\frac{1}{2} \left(-3\pm \sqrt{9-4 j}\right) \]The expression under the radical sign has to be greater than or equal to zero:\[9-4 j\geq 0 \]\[j\leq \frac{9}{4} \]

OpenStudy (anonymous):

Thank you guys so much!

OpenStudy (anonymous):

so should j also be more than 9/4

OpenStudy (anonymous):

i mean\[j \ge -9/4\]

OpenStudy (anonymous):

\[-9/4 \le j \le 9/4\]

OpenStudy (anonymous):

oh i see nevermind it works for all j<9/4 no other condition

OpenStudy (anonymous):

but if j=9/4 only one root twice so shud we conside < or <= @robtobey